# Corona rigidity

### v3

@@ 6.3.3. Countable homogeneity
In the construction of automorphisms, countable saturation can be replaced by a weaker condition. A structure $C$ is countably homogeneous if for every isomorphism $\Phi$ between separable elementary submodels $A$ and $B$ of $C$, a type over $A$ is realized in $C$ if and only if its $\Phi$-image is realized in $C$. Countable homogeneity implies that a structure in a countable language of density $\aleph_{1}$ has $2^{\aleph_{1}}$ automorphisms, by using $\sigma$-complete back-and-forth systems as in §6.1.
By [66, Theorem 1], $\mathcal{Q}(H)$ is not countably homogeneous. Readers not used to continuous logic may find the fact that the offender is a quantifier-free 2-type consisting of one condition amusing. Surprisingly, in [66, Theorem 2] it is shown that even the subgroup of $\mathcal{U}(\mathcal{Q}(H))$ consisting of the unitaries of Fredholm index 0 taken with the metric induced by the norm is not countably homogeneous.
Another failure of countable saturation deserves mention. We will state it in the context of $\mathrm{C}^{*}$-algebras. If $A$ is a separable $\mathrm{C}^{*}$-algebra and $\mathcal{U}$ is a nonprincipal ultrafilter on ${\mathbb{N}}$, then both the (norm) ultrapower $A_{\mathcal{U}}$ and the asymptotic sequence algebra $A_{\infty}=\ell_{\infty}(A)/c_{0}(A)$ are countably saturated (§6.1). Let $\pi_{\mathcal{U}}$ denote the quotient map from $A_{\infty}$ onto $A_{\mathcal{U}}$. A question on the relation between $A_{\infty}$ and $A_{\mathcal{U}}$, motivated by the recent progress on Elliott’s classification programme of $\mathrm{C}^{*}$-algebras, brought the two-sorted structure $\mathfrak{A}=(A_{\infty},A_{\mathcal{U}},\pi_{\mathcal{U}})$ into the focus (see [61]). At the hindsight, it is not too surprising that the countable saturation of $\mathfrak{A}$ is equivalent to the ultrafilter $\mathcal{U}$ being a P-point ([61, Theorem C]).
@@ 6.3.4. Degree-1 saturation
with coefficients $a_{i,j}$ in $C$ and $r\in\mathbb{R}_{+}$. For a degree-1 type $t(\bar{x})$ over $C$ the notions of satisfiablity and realization in $C$ are defined similar to the general types (§6.1), replacing arbitrary conditions with degree-1 conditions. A $\mathrm{C}^{*}$-algebra $C$ is countably degree-1 saturated if every satisfiable countable degree-1 type over $C$ is realized in $C$.
@@
The countable degree-1 saturation has been shown to unify the proofs of several properties of coronas of $\sigma$-unital $\mathrm{C}^{*}$-algebras (see [63]). Infinite-dimensional countably degree-1 saturated $\mathrm{C}^{*}$-algebras have density character at least $\mathfrak{c}$ and possess closure properties resembling those of von Neumann algebras. For instance, if $C$ is a countably degree-1 saturated $\mathrm{C}^{*}$-algebra then there are no $(\aleph_{0},\aleph_{0})$-gaps in $C_{+}$, or in operator algebraic terms $C$ has the countable Riesz separation property (CRISP, for short). Additionally, every countably degree-1 saturated $\mathrm{C}^{*}$-algebra is an SAW-algebra: For any two orthogonal elements $a$ and $b$ in $C_{+}$ there exists $c$ in $C_{+}$ such that $ac=a$ and $bc=0$. The class of SAW-algebras was introduced by Pedersen in [122] as noncommutative analog of sub-Stonean spaces, where it was shown that the corona algebra of any $\sigma$-unital non-unital $\mathrm{C}^{*}$-algebra is an SAW-algebra. SAW-algebras have CRISP ([123]) and they are not isomorphic to a tensor product of two infinite-dimensional $\mathrm{C}^{*}$-algebras ([79]). The latter implies that, for example, the Calkin algebra is not isomorphic to the tensor product of the Calkin algebra with itself, which suggests the existence of (at least to some extent) dimension phenomena associated to the Calkin algebra (see §11.0.1). The class of SAW-algebras is much larger than the class of countably degree-1 saturated algebras since every von Neumann algebra is an SAW-algebra, while no infinite-dimensional von Neumann algebra is countably degree-1 saturated. For a more complete list of properties of countably degree-1 saturated $\mathrm{C}^{*}$-algebras, see [59, §15].
@@ 6.4. Stratifying $\mathcal{Q}(H)$
An easy diagonalization argument shows that $(\operatorname{Part_{{\mathbb{N}}}},\leq^{*})$ is $\sigma$-directed (for the aficionados of Tukey ordering, $(\operatorname{Part_{{\mathbb{N}}}},\leq^{*})$ is cofinally equivalent to $({\mathbb{N}}^{\mathbb{N}},\leq^{*})$ by [59, Theorem 9.7.8]).
@@ 6.5. Cohomology and automorphisms
At the first glance, the outline of the original construction of an outer automorphism of the Calkin algebra ([124]) using $\mathrm{CH}$ is unsurprising. The automorphism is constructed recursively in $\aleph_{1}$ steps, assuring that its restriction to every separable $\mathrm{C}^{*}$-subalgebra is implemented by a unitary. (In fact, the construction produces $2^{\aleph_{1}}$ distinct outer automorphisms, but this is not surprising either.) At a closer glance, this is one of the subtlest $\mathrm{CH}$ constructions known to the authors. As could be expected, the subtlety of this approach shows at the limit stages. At the first limit stage, one has an increasing sequence $A_{n}$, for $n\in{\mathbb{N}}$, of separable $\mathrm{C}^{*}$-subalgebras of $\mathcal{Q}(H)$ and unitaries $u_{n}$ such that for $m<n$ the actions of $u_{m}$ and $u_{n}$ on $A_{m}$ agree. In order to continue the construction, one needs to choose a single unitary $u$ that implements the action of $u_{n}$ on $A_{n}$ for all $n$. The construction of two non-unitarily equivalent unital copies of the CAR algebra in Example 6.16 and Example 6.17 shows that this is in general impossible. The latter example gives a sequence of unitaries $u_{n}$ whose actions on a unital copy of the CAR algebra in $\mathcal{Q}(H)$ cohere, but the Fredholm indices of $u_{n}$ converge to $\infty$ and $u_{n}$ cannot be uniformized by a single unitary. In [124], the unitaries implementing partial isomorphisms are assumed to be homotopic. The construction of a unitary extending the paths constructed up to that point uses difficult ideas developed in relation to Kasparov’s KK theory.
A simpler—to a logician—construction of an outer automorphism of $\mathcal{Q}(H)$ was described in [58, §1] and extended to coronas of some other $\sigma$-unital $\mathrm{C}^{*}$-algebras in [29] (see [59, §17.1]).
@@
The proof relies on §6.4. Let $\mathbb{T}$ denote the multiplicative group of complex numbers of modulus 1. Identify $\mathbb{T}^{\mathbb{N}}$ with the unitary group of $\ell_{\infty}({\mathbb{N}})$, and therefore with the group of unitaries in $\mathcal{B}(H)$ diagonalized by a distinguished basis of $H$.
@@ 6.5.1. An inverse limit
In [59], all coherent families are by definition total. Our modification of terminology was made in order to make it possible to say that an easy diagonalization argument shows that every countable coherent family of unitaries is trivial.
@@ 6.6. Higher-dimensional Čech–Stone remainders
So far we discussed how to construct topologically nontrivial automorphisms for coronas which are countably saturated (§6.1), or coronas of stabilizations of unital separable $\mathrm{C}^{*}$-algebras (§6.5).
There are though plenty of algebras to which the arguments we exposed so far do not apply. A good example is $C_{0}(\mathbb{R}^{2})$. [^22] The corona $C_{0}(\mathbb{R})$ is countably saturated (see [71]) and therefore it has $2^{\mathfrak{c}}$ automorphisms that are not topologically trivial under $\mathrm{CH}$. These were also constructed, via topological methods, by Yu, see [85]. Its corona is not saturated ([74, Corollary 2]) and, being abelian, its only inner automorphism is the identity, making the methods of §6.5 completely ineffective.
@@
The $D_{\alpha}$ consists of functions which are more and more Lipschitz when restricted to $X_{n}$ (see for details [169, §2]).
@@ 6.7. Other models for non-rigidity
Finally, by a result of Baumgartner reproduced in [167, p. 130], the existence of an $\aleph_{1}$-generated P-point ultrafilter $\mathcal{U}$ implies the existence of a nontrivial automorphism of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ that fixes $\mathcal{U}$. On the other hand, an automorphism of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ is nowhere trivial if its restriction to $\mathcal{P}(A)/\operatorname{Fin}$ is not trivial for any infinite $A\subseteq{\mathbb{N}}$ (the restriction of Baumgartner’s automorphism to any set in $\mathcal{U}$ is trivial). It is possible to construct such automorphisms from CH, and in [150] they were constructed from an assumption reminiscent of $\mathfrak{d}=\aleph_{1}$.
@@ 7.1. Shelah’s proof
This proof (first appearing in [141, §IV], then in [144, §IV], and in an improved form, allowing $\mathfrak{c}$ to be arbitrarily large, in [93]) started the entire subject that this survey is about. Since several features of Shelah’s proof still, over forty years later, resonate throughout the subject, we take the opportunity to sketch it.
@@
If this recursive construction can be completed, then $\Phi$ is ‘nipped in the bud’ by $\mathbb{T}$, as desired. We proceed to analyze the situation in which the construction halts at stage $\alpha$. Assume for a moment that $B_{\alpha}$ has been fixed. For $A\subseteq B_{\alpha}$ write $\mathbb{T}[A]$ for $\mathbb{T}_{(\bar{B}^{\frown}B_{\alpha},\bar{A}^{\frown}A)}$. If (7.1.2) applies, then the set
@@
From $q$ as in (7.1) one can extract a partial function $g_{q}$ whose restriction to $B_{\alpha}$ is equal to $g$ modulo finite (let’s say that $g_{q}$ ‘works’ on $B_{\alpha}$). This is true for every eligible candidate for $B_{\alpha}$. Since (7.1) does not refer to $B_{\alpha}$, there is a countable set of functions, $g_{q}$, for $q\in\mathbb{T}^{\prime}$, that between them serve as a lift of $\Phi$ on $\mathcal{P}(B)$, for every $B\in\{B_{\beta}\mid\beta<\alpha\}^{\perp}\cap{\mathcal{J}}_{\Phi}$. The ideal =${\mathcal{I}}=\{B_{\beta}\mid\beta<\alpha\}^{\perp}$ is included in ${\mathcal{J}}_{\Phi}$. [^27] This is the white lie promised in footnote 25. As an orthogonal of a countable set, it is countably directed under $\subseteq^{*}$, and therefore a single function $g_{q}$ ‘works’ on a cofinal set of $B\in{\mathcal{I}}$. An easy diagonalization argument shows that there is $k\in{\mathbb{N}}$ such that $g_{q}$ ‘works’ on ${\mathbb{N}}\setminus\bigcup_{j<k}B_{j}$, hence this set belongs to ${\mathcal{J}}_{\Phi}$. Thus ${\mathbb{N}}$, as the union of $k+1$ sets in ${\mathcal{J}}_{\Phi}$, belongs to ${\mathcal{J}}_{\Phi}$ and $\Phi$ is trivial; contradiction.
@@
By using a modification of the oracle-c.c., Just proved that the assertion that all automorphisms of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ are trivial is consistent with the continuum being arbitrarily large ([93, Theorem A]). More importantly for our story, he also proved that in his model $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ cannot be embedded into $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ for any Borel ideal that is not generated by a single set over $\operatorname{Fin}$ ([93, Theorem 0.6]). [^29] Just’s proof used a measurable cardinal, but this assumption has later been removed. Just proved that in the same model every isomorphism between quotients over Borel ideals is topologically trivial on a large set (see [92] for a nice application).
@@ 7.2. Quotients of $\mathcal{P}({\mathbb{N}})$ and zero-dimensional remainders
For a set ${\mathcal{X}}$, let $[{\mathcal{X}}]^{2}$ denote the set of unordered pairs of elements of ${\mathcal{X}}$. If ${\mathcal{X}}$ is a topological space, we say that $L\subseteq[{\mathcal{X}}]^{2}$ is open if $\{(x,y)\in{\mathcal{X}}^{2}\mid\{x,y\}\in L\}$ is open. The following was isolated and proved to follow from $\mathrm{PFA}$ in [160, §8].
@@
$\mathrm{OCA_{T}}$ asserts that for every separably metrizable space ${\mathcal{X}}$ and an open $L\subseteq[{\mathcal{X}}]^{2}$ either there exists an uncountable ${\mathcal{Y}}\subseteq{\mathcal{X}}$ such that $[{\mathcal{Y}}]^{2}\subseteq L$ or there is a partition ${\mathcal{X}}=\bigcup_{n}{\mathcal{X}}_{n}$ such that each $[{\mathcal{X}}_{n}]^{2}$ is disjoint from $L$.
@@
The following is an initial segment of [166, Lemma 2.2].
@@
In the following step of the proof that $\Phi$ is topologically trivial one shows that if $\Phi$ has a $\sigma$-Borel lifting on $\mathcal{P}(A)$ and $A=\bigcup_{n}A_{n}$ then $\Phi$ has a Borel lifting (i.e., is topologically trivial) on some $\mathcal{P}(A_{n})$. An elegant proof of this using the Kuratowski–Ulam Theorem is due to Fremlin, given in a more general context in [50, Lemma 3.12.3]. This, together with Lemma 7.3 and $\mathrm{MA}$, shows that the ideal ${\mathcal{J}}_{\Phi}$ of $A\subseteq{\mathbb{N}}$ such that $\Phi$ is topologically trivial on $\mathcal{P}(A)$ is large (e.g., it is nonmeager, but the details are not relevant for our purposes).
@@
The following is a special case of [94, Theorem 11] (Just considered $\Phi$ which is a lattice embedding and obtained a slightly weaker conclusion).
@@
Solecki’s characterization of analytic P-ideals (Theorem 3.6) associates a sequence of closed approximations to every analytic P-ideal by Lemma 7.6. Together with a meticulous uniformization of locally defined liftings ([50, §3.13]) one obtains the following (see [50, Theorem 3.3.5]).
@@
Assume $\mathrm{OCA_{T}}$ and $\mathrm{MA}_{\aleph_{1}}$ and let $\Phi\colon\mathcal{P}({\mathbb{N}})\to\mathcal{P}({\mathbb{N}})/{\mathcal{J}}$ be a homomorphism such that ${\mathcal{J}}\supseteq\operatorname{Fin}$ is countably generated or an analytic P-ideal. Then $\Phi$ is the direct sum of a topologically trivial homomorphisms and one with a nonmeager kernel. In particular, every isomorphism between a quotient over an analytic ideal and $\mathcal{P}({\mathbb{N}})/{\mathcal{J}}$ is topologically trivial. [^32] It is well-known that every proper analytic ideal that includes $\operatorname{Fin}$ is meager.
@@
Assume $\mathrm{OCA_{T}}$ and $\mathrm{MA}_{\aleph_{1}}$. Let $\mathcal{I}$ and $\mathcal{J}$ be Borel ideals in $\mathcal{P}({\mathbb{N}})$ such that $\mathcal{J}$ be either a nonpathological $P$-ideal or countably generated. Then
@@
This is a good moment to state [55, Question 15.2] (3204 is a poor estimate, but not a joke, see [55, Theorem 10.4]).
@@
Suppose that ${\mathcal{I}}$ is an $F_{\sigma\delta}$ ideal on ${\mathbb{N}}$. Are there closed approximations $\mathcal{K}_{m}$, for $m\in{\mathbb{N}}$, such that ${\mathcal{I}}=\bigcap_{m}(\mathcal{K}_{m}^{d}\underline{\cup}\operatorname{Fin})$ for every $d\geq 1$? What about the cases $d=1$ and $d=3204$?
An ideal for which the answer to Question 7.10 is positive is said to be 3204-determined by closed approximations. By [55, Theorem 10.4], $\mathrm{PFA}$ implies that for such ${\mathcal{I}}$ and any analytic ideal ${\mathcal{J}}$, every isomorphism between $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ and $\mathcal{P}({\mathbb{N}})/{\mathcal{J}}$ is topologically trivial. This class includes the ideals $\operatorname{NWD}(\mathbb{Q})$, $\operatorname{NULL}(\mathbb{Q})$, and the Weyl ideal ${\mathcal{Z}}_{W}$ (see Theorem 5.13).

### v4

@@ 6.3.3. Countable homogeneity
In the construction of automorphisms, countable saturation can be replaced by a weaker condition. A structure $C$ is countably homogeneous if for every isomorphism $\Phi$ between separable elementary submodels $A$ and $B$ of $C$, a type over $A$ is realised in $C$ if and only if its $\Phi$-image is realised in $C$. Countable homogeneity implies that a structure in a countable language of density $\aleph_{1}$ has $2^{\aleph_{1}}$ automorphisms, by using $\sigma$-complete back-and-forth systems as in §6.1.
By [70, Theorem 1], $\mathcal{Q}(H)$ is not countably homogeneous. Readers not used to continuous logic may find the fact that the offender is a quantifier-free 2-type consisting of one condition amusing. Surprisingly, in [70, Theorem 2] it is shown that even the subgroup of $\mathcal{U}(\mathcal{Q}(H))$ consisting of the unitaries of Fredholm index 0 taken with the metric induced by the norm is not countably homogeneous.
Another failure of countable saturation deserves mention. We will state it in the context of $\mathrm{C}^{*}$-algebras. If $A$ is a separable $\mathrm{C}^{*}$-algebra and $\mathcal{U}$ is a nonprincipal ultrafilter on ${\mathbb{N}}$, then both the (norm) ultrapower $A_{\mathcal{U}}$ and the asymptotic sequence algebra $A_{\infty}=\ell_{\infty}(A)/c_{0}(A)$ are countably saturated (§6.1). Let $\pi_{\mathcal{U}}$ denote the quotient map from $A_{\infty}$ onto $A_{\mathcal{U}}$. A question on the relation between $A_{\infty}$ and $A_{\mathcal{U}}$, motivated by the recent progress on Elliott’s classification programme of $\mathrm{C}^{*}$-algebras, brought the two-sorted structure $\mathfrak{A}=(A_{\infty},A_{\mathcal{U}},\pi_{\mathcal{U}})$ into the focus (see [63]). At the hindsight, it is not too surprising that the countable saturation of $\mathfrak{A}$ is equivalent to the ultrafilter $\mathcal{U}$ being a P-point ([63, Theorem C]).
@@ 6.3.4. Degree-1 saturation
with coefficients $a_{i,j}$ in $C$ and $r\in\mathbb{R}_{+}$. For a degree-1 type $t(\bar{x})$ over $C$ the notions of satisfiablity and realisation in $C$ are defined similar to the general types (§6.1), replacing arbitrary conditions with degree-1 conditions. A $\mathrm{C}^{*}$-algebra $C$ is countably degree-1 saturated if every satisfiable countable degree-1 type over $C$ is realised in $C$.
@@
The countable degree-1 saturation has been shown to unify the proofs of several properties of coronas of $\sigma$-unital $\mathrm{C}^{*}$-algebras (see [67]). Infinite-dimensional countably degree-1 saturated $\mathrm{C}^{*}$-algebras have density character at least $\mathfrak{c}$ and possess closure properties resembling those of von Neumann algebras. For instance, if $C$ is a countably degree-1 saturated $\mathrm{C}^{*}$-algebra then there are no $(\aleph_{0},\aleph_{0})$-gaps in $C_{+}$, or in operator algebraic terms $C$ has the countable Riesz separation property (CRISP, for short). Additionally, every countably degree-1 saturated $\mathrm{C}^{*}$-algebra is an SAW-algebra: For any two orthogonal elements $a$ and $b$ in $C_{+}$ there exists $c$ in $C_{+}$ such that $ac=a$ and $bc=0$. The class of SAW-algebras was introduced by Pedersen in [128] as noncommutative analog of sub-Stonean spaces, where it was shown that the corona algebra of any $\sigma$-unital non-unital $\mathrm{C}^{*}$-algebra is an SAW-algebra. SAW-algebras have CRISP ([129]) and they are not isomorphic to a tensor product of two infinite-dimensional $\mathrm{C}^{*}$-algebras ([83]). The latter implies that, for example, the Calkin algebra is not isomorphic to the tensor product of the Calkin algebra with itself, which suggests the existence of (at least to some extent) dimension phenomena associated to the Calkin algebra (see §11.0.1). The class of SAW-algebras is much larger than the class of countably degree-1 saturated algebras since every von Neumann algebra is an SAW-algebra, while no infinite-dimensional von Neumann algebra is countably degree-1 saturated. For a more complete list of properties of countably degree-1 saturated $\mathrm{C}^{*}$-algebras, see [61, §15].
@@ 6.4. Stratifying $\mathcal{Q}(H)$
An easy diagonalisation argument shows that $(\operatorname{Part_{{\mathbb{N}}}},\leq^{*})$ is $\sigma$-directed (for the aficionados of Tukey ordering, $(\operatorname{Part_{{\mathbb{N}}}},\leq^{*})$ is cofinally equivalent to $({\mathbb{N}}^{\mathbb{N}},\leq^{*})$ by [61, Theorem 9.7.8]).
@@ 6.5. Cohomology and automorphisms
At the first glance, the outline of the original construction of an outer automorphism of the Calkin algebra ([130]) using $\mathrm{CH}$ is unsurprising. The automorphism is constructed recursively in $\aleph_{1}$ steps, assuring that its restriction to every separable $\mathrm{C}^{*}$-subalgebra is implemented by a unitary. (In fact, the construction produces $2^{\aleph_{1}}$ distinct outer automorphisms, but this is not surprising either.) At a closer glance, this is one of the subtlest $\mathrm{CH}$ constructions known to the authors. As could be expected, the subtlety of this approach shows at the limit stages. At the first limit stage, one has an increasing sequence $A_{n}$, for $n\in{\mathbb{N}}$, of separable $\mathrm{C}^{*}$-subalgebras of $\mathcal{Q}(H)$ and unitaries $u_{n}$ such that for $m<n$ the actions of $u_{m}$ and $u_{n}$ on $A_{m}$ agree. In order to continue the construction, one needs to choose a single unitary $u$ that implements the action of $u_{n}$ on $A_{n}$ for all $n$. The construction of two non-unitarily equivalent unital copies of the CAR algebra in Example 6.16 and Example 6.17 shows that this is in general impossible. The latter example gives a sequence of unitaries $u_{n}$ whose actions on a unital copy of the CAR algebra in $\mathcal{Q}(H)$ cohere, but the Fredholm indices of $u_{n}$ converge to $\infty$ and $u_{n}$ cannot be uniformised by a single unitary. In [130], the unitaries implementing partial isomorphisms are assumed to be homotopic. The construction of a unitary extending the paths constructed up to that point uses difficult ideas developed in relation to Kasparov’s KK theory.
A simpler—to a logician—construction of an outer automorphism of $\mathcal{Q}(H)$ was described in [60, §1] and extended to coronas of some other $\sigma$-unital $\mathrm{C}^{*}$-algebras in [29] (see [61, §17.1]).
@@
The proof relies on §6.4. Let $\mathbb{T}$ denote the multiplicative group of complex numbers of modulus 1. Identify $\mathbb{T}^{\mathbb{N}}$ with the unitary group of $\ell_{\infty}({\mathbb{N}})$, and therefore with the group of unitaries in $\mathcal{B}(H)$ diagonalised by a distinguished basis of $H$.
@@ 6.5.1. An inverse limit
In [61], all coherent families are by definition total. Our modification of terminology was made in order to make it possible to say that an easy diagonalisation argument shows that every countable coherent family of unitaries is trivial.
@@ 6.6. Higher-dimensional Čech–Stone remainders
So far we discussed how to construct topologically nontrivial automorphisms for coronas which are countably saturated (§6.1), or coronas of stabilisations of unital separable $\mathrm{C}^{*}$-algebras (§6.5).
There are though plenty of algebras to which the arguments we exposed so far do not apply. A good example is $C_{0}(\mathbb{R}^{2})$. [^22] The corona $C_{0}(\mathbb{R})$ is countably saturated (see [75]) and therefore it has $2^{\mathfrak{c}}$ automorphisms that are not topologically trivial under $\mathrm{CH}$. These were also constructed, via topological methods, by Yu, see [89]. Its corona is not saturated ([78, Corollary 2]) and, being abelian, its only inner automorphism is the identity, making the methods of §6.5 completely ineffective.
@@
The $D_{\alpha}$ consists of functions which are more and more Lipschitz when restricted to $X_{n}$ (see for details [177, §2]).
@@ 6.7. Other models for non-rigidity
Finally, by a result of Baumgartner reproduced in [174, p. 130], the existence of an $\aleph_{1}$-generated P-point ultrafilter $\mathcal{U}$ implies the existence of a nontrivial automorphism of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ that fixes $\mathcal{U}$. On the other hand, an automorphism of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ is nowhere trivial if its restriction to $\mathcal{P}(A)/\operatorname{Fin}$ is not trivial for any infinite $A\subseteq{\mathbb{N}}$ (the restriction of Baumgartner’s automorphism to any set in $\mathcal{U}$ is trivial). It is possible to construct such automorphisms from CH, and in [156] (see also [158]) they were constructed from an assumption reminiscent of $\mathfrak{d}=\aleph_{1}$.
@@ 7.1. Shelah’s proof
This proof (first appearing in [147, §IV], then in [150, §IV], and in an improved form, allowing $\mathfrak{c}$ to be arbitrarily large, in [97]) started the entire subject that this survey is about. Since several features of Shelah’s proof still, over forty years later, resonate throughout the subject, we take the opportunity to sketch it.
@@
If this recursive construction can be completed, then $\Phi$ is ‘nipped in the bud’ by $\mathbb{T}$, as desired. We proceed to analyse the situation in which the construction halts at stage $\alpha$. Assume for a moment that $B_{\alpha}$ has been fixed. For $A\subseteq B_{\alpha}$ write $\mathbb{T}[A]$ for $\mathbb{T}_{(\bar{B}^{\frown}B_{\alpha},\bar{A}^{\frown}A)}$. If (7.1.2) applies, then the set
@@
From $q$ as in (7.1) one can extract a partial function $g_{q}$ whose restriction to $B_{\alpha}$ is equal to $g$ modulo finite (let’s say that $g_{q}$ ‘works’ on $B_{\alpha}$). This is true for every eligible candidate for $B_{\alpha}$. Since (7.1) does not refer to $B_{\alpha}$, there is a countable set of functions, $g_{q}$, for $q\in\mathbb{T}^{\prime}$, that between them serve as a lift of $\Phi$ on $\mathcal{P}(B)$, for every $B\in\{B_{\beta}\mid\beta<\alpha\}^{\perp}\cap{\mathcal{J}}_{\Phi}$. The ideal =${\mathcal{I}}=\{B_{\beta}\mid\beta<\alpha\}^{\perp}$ is included in ${\mathcal{J}}_{\Phi}$. [^27] This is the white lie promised in footnote 25. As an orthogonal of a countable set, it is countably directed under $\subseteq^{*}$, and therefore a single function $g_{q}$ ‘works’ on a cofinal set of $B\in{\mathcal{I}}$. An easy diagonalisation argument shows that there is $k\in{\mathbb{N}}$ such that $g_{q}$ ‘works’ on ${\mathbb{N}}\setminus\bigcup_{j<k}B_{j}$, hence this set belongs to ${\mathcal{J}}_{\Phi}$. Thus ${\mathbb{N}}$, as the union of $k+1$ sets in ${\mathcal{J}}_{\Phi}$, belongs to ${\mathcal{J}}_{\Phi}$ and $\Phi$ is trivial; contradiction.
@@
By using a modification of the oracle-c.c., Just proved that the assertion that all automorphisms of $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ are trivial is consistent with the continuum being arbitrarily large ([97, Theorem A]). More importantly for our story, he also proved that in his model $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ cannot be embedded into $\mathcal{P}({\mathbb{N}})/\operatorname{Fin}$ for any Borel ideal that is not generated by a single set over $\operatorname{Fin}$ ([97, Theorem 0.6]). [^29] Just’s proof used a measurable cardinal, but this assumption has later been removed. Just proved that in the same model every isomorphism between quotients over Borel ideals is topologically trivial on a large set (see [96] for a nice application).
@@ 7.2. Quotients of $\mathcal{P}({\mathbb{N}})$ and zero-dimensional remainders
For a set ${\mathcal{X}}$, let $[{\mathcal{X}}]^{2}$ denote the set of unordered pairs of elements of ${\mathcal{X}}$. If ${\mathcal{X}}$ is a topological space, we say that $L\subseteq[{\mathcal{X}}]^{2}$ is open if $\{(x,y)\in{\mathcal{X}}^{2}\mid\{x,y\}\in L\}$ is open. The following was isolated and proved to follow from $\mathrm{PFA}$ in [167, §8].
@@
$\mathrm{OCA_{T}}$ asserts that for every separably metrisable space ${\mathcal{X}}$ and an open $L\subseteq[{\mathcal{X}}]^{2}$ either there exists an uncountable ${\mathcal{Y}}\subseteq{\mathcal{X}}$ such that $[{\mathcal{Y}}]^{2}\subseteq L$ or there is a partition ${\mathcal{X}}=\bigcup_{n}{\mathcal{X}}_{n}$ such that each $[{\mathcal{X}}_{n}]^{2}$ is disjoint from $L$.
@@
The following is an initial segment of [173, Lemma 2.2].
@@
In the following step of the proof that $\Phi$ is topologically trivial one shows that if $\Phi$ has a $\sigma$-Borel lifting on $\mathcal{P}(A)$ and $A=\bigcup_{n}A_{n}$ then $\Phi$ has a Borel lifting (i.e., is topologically trivial) on some $\mathcal{P}(A_{n})$. An elegant proof of this using the Kuratowski–Ulam Theorem is due to Fremlin, given in a more general context in [52, Lemma 3.12.3]. This, together with Lemma 7.3 and $\mathrm{MA}$, shows that the ideal ${\mathcal{J}}_{\Phi}$ of $A\subseteq{\mathbb{N}}$ such that $\Phi$ is topologically trivial on $\mathcal{P}(A)$ is large (e.g., it is nonmeager, but the details are not relevant for our purposes).
@@
The following is a special case of [98, Theorem 11] (Just considered $\Phi$ which is a lattice embedding and obtained a slightly weaker conclusion).
@@
Solecki’s characterisation of analytic P-ideals (Theorem 3.6) associates a sequence of closed approximations to every analytic P-ideal by Lemma 7.6. Together with a meticulous uniformisation of locally defined liftings ([52, §3.13]) one obtains the following (see [52, Theorem 3.3.5]).
@@
Assume $\mathrm{OCA_{T}}$ and $\mathrm{MA}$ and let $\Phi\colon\mathcal{P}({\mathbb{N}})\to\mathcal{P}({\mathbb{N}})/{\mathcal{J}}$ be a homomorphism such that ${\mathcal{J}}\supseteq\operatorname{Fin}$ is countably generated or an analytic P-ideal. Then $\Phi$ is the direct sum of a topologically trivial homomorphisms and one with a nonmeager kernel. In particular, every isomorphism between a quotient over an analytic ideal and $\mathcal{P}({\mathbb{N}})/{\mathcal{J}}$ is topologically trivial. [^32] It is well-known that every proper analytic ideal that includes $\operatorname{Fin}$ is meager.
@@
Assume $\mathrm{OCA_{T}}$ and $\mathrm{MA}$. Let $\mathcal{I}$ and $\mathcal{J}$ be Borel ideals in $\mathcal{P}({\mathbb{N}})$ such that $\mathcal{J}$ be either a nonpathological $P$-ideal or countably generated. Then
@@
This is a good moment to state a revised version of [57, Question 15.2] (the original question asked for $d=3204$ in place of $d=80$; compare [64, Theorem 1.4] and [57, Theorem 10.4]).
@@
Suppose that ${\mathcal{I}}$ is an $F_{\sigma\delta}$ ideal on ${\mathbb{N}}$. Are there closed approximations $\mathcal{K}_{m}$, for $m\in{\mathbb{N}}$, such that ${\mathcal{I}}=\bigcap_{m}(\mathcal{K}_{m}^{d}\underline{\cup}\operatorname{Fin})$ for every $d\geq 1$? What about the cases $d=1$ and $d=80$?
An ideal for which the answer to Question 7.10 is positive is said to be $80$-determined by closed approximations. By [64, Theorem 1.4] (see [57, Theorem 10.4]), $\mathrm{OCA_{T}}$ and $\mathrm{MA_{\sigma-linked}}$ together imply that for such ${\mathcal{I}}$ and any analytic ideal ${\mathcal{J}}$, every isomorphism between $\mathcal{P}({\mathbb{N}})/{\mathcal{I}}$ and $\mathcal{P}({\mathbb{N}})/{\mathcal{J}}$ is topologically trivial. This class includes the ideals $\operatorname{NWD}(\mathbb{Q})$, $\operatorname{NULL}(\mathbb{Q})$, and the Weyl ideal ${\mathcal{Z}}_{W}$ (see Theorem 5.13).